The masterfoods company manufactures bags of peanut butter M&Ms. they report that they make 10% each brown and red candies, and 20% each yellow, blue, and orange candies. The rest of the candies are green.

If you pick four M&Ms in a row, what is the probability that they are:

A all blue?
B none are green?
C at least one is red?
D the fourth one is the first one that is brown?
Answer #1

Probability of brown candies = 0.10

Probability of red candies = 0.10

Probability of yellow candies = 0.20

Probability of blue candies = 0.20

Probability of orange candies = 0.20

Probability of green candies =1-0.10-0.10-0.20-0.20-0.20 = 0.20

a) Probability of all blue candies, P(Blue) = (0.20)4 = 0.0016

b) Probability that none is green, P(None green) = (0.80)4 =  0.4096

c) Probability of at least one red, P(at least one red) = 1 - P(No red)

= 1 - (0.90)4 = 1 - 0.6561 = 0.3439

d) Probability that the fourth one is the first one that is brown = P(not brown)*P(not brown)*P(not brown)*P(brown)

= 0.90*0.90*0.90*0.10 = 0.0729